OpenStudy (anonymous):

Write the slope-intercept form of an equation of the line passing through the points (1, 2) and (5, -10). Show your work

OpenStudy (anonymous):

y = mx + b is slope intercept so first have to find the slope using the equation y2-y1 = x2-1 and plug that value in for m and solve for b

OpenStudy (anonymous):

Pick any point for x from the first or second equation

OpenStudy (anonymous):

I would go with the first one